OpenStudy (anonymous):

Solve. 16b^2 -25=0 SHOW STEPS PLEASE

OpenStudy (alexandervonhumboldt2):

a^2-b^2=(a-b)*(a+b) 4^2b^2-5^2=0 (4b-5)*(4b+5)=0 b=5/4 b=-5/4